OpenStudy (elise_a18):

3(x + 1) = -2(x - 1) + 6 Solve for X

OpenStudy (elise_a18):

I got 5 but I think that's incorrect

Mehek (mehek14):

use distributive property first

Mehek (mehek14):

*first*

OpenStudy (elise_a18):

i did

Mehek (mehek14):

ok then let me check your answer

OpenStudy (elise_a18):

5 ._.

Mehek (mehek14):

3(5 + 1) = -2(5 - 1) + 6 3(6) = -2(4) + 6 18 = -8 + 6 18 = -2 not correct

OpenStudy (elise_a18):

k...

Mehek (mehek14):